# Discrete Diffusion-Based Model-Level Explanation of Heterogeneous GNNs with Node Features

### v1

@@ 2. Node feature generation
We conduct feature selection experiments for varying numbers of selected features (2, 4, 8, 50 for DBLP and 2, 5, 10, 3066 for IMDB) using Variance Threshold (see Table 2) and top $k$ frequent feature selection methods. For both feature selection methods, we obtain good explanation graphs for a large feature size of 50 using DiTabDDPM for DBLP.
Runtime Analysis
@@
Unlike graph generation techniques in general, which evaluate performance in terms of scalability as the graph size increases, we evaluate the scalability of our approach as the size of the original dataset increases. We mainly compare our runtime with the existing diffusion-based approach D4Explainer (Table 4). D4Explainer is faster than DiGNNExplainer for small synthetic datasets like Tree-Cycle (871 nodes), but the runtime increases with increasing dataset size for IMDB (11,616 nodes) and DBLP (26,128 nodes). We take into account the time needed for all graph sizes. In comparison, the runtime of our approach is constant for both small and large datasets.
@@
Table 4: Runtime (in secs.) comparisons with D4Explainer baseline for all datasets.
@@
D4Explainer 39174.97 19406.46 658.26 5702.66 658.59 821.59 800.71
DiGNNExplainer 5895.05 5006.71 6910.2 4637.09 5191.56 5500.54 1992.29
Case Study
Case study examples of 1) Explanation graphs with node features and 2) Explanation graphs generated for the same class for each approach, i.e., Class 1, along with example motifs for each dataset can be found in the appendix.
Discussion
@@
We discuss the following improvements and extensions to our approach: (1) Our work separates the process of generating explanation graphs into two distinct steps: graph structure generation and node feature generation, which use separate models. The advantage of this methodology is that we generate realistic features, i.e., we learn features using specific techniques, discrete/continuous, using different models instead of a single model as in VAE. The disadvantage is that the potential dependence of node features on the graph structure is not captured. We address this dependence in DiTabDDPM and TabDDPM, but the generated features show a reduced cosine similarity. Hence, we show our main results without considering graph structure dependence. (2) We obtain consistent runtime for small and large datasets, which is crucial for practical applications since real-world datasets are large. In future work, it can be made faster by using multiple GPUs simultaneously, as done by the recent code of DiGress. (3) We train a separate DiGress model for each graph size which may introduce significant computational costs, but our approach is more robust. This is because, while DiGress allows to input and sample multiple graph sizes using a single diffusion model, the diffusion training process considers all graphs to be of the highest node size, which is 15 for DBLP. Our ablation studies for training multiple sizes using a single model show comparatively poor evaluation results and generate a higher number of disconnected graphs. (4) Our approach is easily extendable to real-world heterogeneous directed graphs, since our GNNs (GraphSAGE and HAN) can be trained on directed graphs. Moreover, DiGNNExplainer, which builds upon DiGress, is able to support directed graphs with only minor modifications, i.e., by modifying the way noise is added to the adjacency matrix for directed graphs. (5) We obtain one explanation for each class, but our current implementation also supports multiple explanation graphs, i.e., the top $k$ explanations for each class, sorted by softmax probabilities.
Conclusion
@@
Our approach, DiGNNExplainer, is the first to explain heterogeneous GNNs on a model-level for real-world datasets using explanation graphs with actual node features. Our experimental results demonstrate that incorporating actual node features significantly improves explanation faithfulness. Specifically, DiGNNExplainer achieves higher predictive and ground-truth faithfulness compared to state-of-the-art baselines, while also producing explanation graphs that closely match the statistical properties of real-world graphs. Key aspects of our approach that contribute to the generation of plausible and faithful explanations include: (1) Node feature generation via our novel discrete diffusion model, DiTabDDPM, which enables the synthesis of realistic discrete features; and (2) our diffusion-based approach for modeling the underlying heterogeneous relations of graphs, supporting the generation of structurally realistic explanation graphs. Beyond the citation and movie datasets used in our experiments, DiGNNExplainer is broadly applicable to any domain with heterogeneous graph data and structured node features—such as healthcare, finance, or chemistry. Our main goal was to support explanation generation with actual node features to improve the explainability of heterogeneous GNNs. Further runtime optimization is left for future work.
@@ Appendix A Diffusion Models
Figure 2: Diffusion steps for generation of synthetic graph and discrete node features.
DiGress
DiGress (Vignac et al. 2023) is a discrete diffusion approach for graph generation, which adds noise through discrete time steps based on a Markov process. The noise model is a Markov process over a discrete state-space. In a discrete state-space, each node and edge transitions to a different state with the addition of noise at every time step. A state refers to a node type or an edge type. Following each transition, the node or edge changes to another type. The probabilities of transition from one state to another are represented as transition matrices for nodes and for edges. In DiGress, the next step $G_{t}$ (Figure 2) is obtained by the product of the transition matrix, $Q_{t}$ at step $t$ and the graph at the previous step, $G_{t-1}$. The transition probability, $q$ at $G_{t}$ is given as:
$q(G_{t}|G_{t-1})=G_{t-1}Q_{t}$ (1)
@@
where node loss, $L_{Nodes}$ and edge loss, $L_{Edges}$ represent the cross entropy loss between the true node and edge, and predicted probabilities of each node and edge, and $\lambda\in\mathbb{R}^{+}$ assigns relative importance to nodes and edges. The graph loss is computed as a combined loss of individual node loss and edge loss. The loss of each node and edge is measured using cross entropy. DiGress models node features along with the graph structure. The node features are one-hot encoding of node types. Extra features are added to nodes, including structural information that encodes the presence of cycles, and spectral features for positional encodings needed for a transformer model. These extra features are added at each diffusion step, as the message passing networks are unable to capture the presence of cycles. DiGress can be used to generate small heterogeneous graphs with the same node size as the input graph. Discrete node types and edge types are sampled from the complete noise distributions. The samples are input to a denoising network, which uses self-attention to predict the noise and outputs the distributions for node types and edge types for each previous step.
TabDDPM
TabDDPM (Kotelnikov et al. 2023) extends the denoising diffusion probabilistic (DDPM) approach used for images to synthesize tabular data. In TabDDPM, an input feature vector, $x=[x_{num},x_{cat_{1}}\dots x_{cat_{C}}]$ consists of both numeric features, $x_{num}$ and categorical features, $x_{cat_{i}}$ with $i\in C$ categories. The categorical features are represented as one-hot encoding of categories. TabDDPM models categorical and binary features using multinomial diffusion and numerical features using continuous noise. The forward process adds Gaussian noise using a Markov process, while the reverse process uses a denoising neural network consisting of MLP layers with SiLU non-linear activation. The denoising network is an MLP architecture of the form (Gorishniy et al. 2021):
$\mathit{MLP}(x)=Linear(\mathit{MLPBlock}(\dots(\mathit{MLPBlock}(x)))$ (2)
@@ Appendix B Additional Details for DiGNNExplainer
DiTabDDPM - Denoising network
@@
Node Feature Reconstruction Loss
@@
$L_{recon}=\mathbb{E}[log_{p_{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{t})]$
Postprocessing
Two nodes $\tau_{1},\tau_{2}$ in the metagraph are connected if and only if there exists two nodes $v_{1}$ and $v_{2}$ in the graph such that $v_{1}$ has type $\tau_{1}$ and $v_{2}$ has type $\tau_{2}$. The generated graph is consistent if and only if the metagraph of the generated graph is a subgraph of the metagraph of the dataset.
Explanation Graph Selection
@@ Appendix C Datasets
Table 5: Dataset Statistics: number of graphs, average number of nodes per graph, number of classified nodes (e.g., author nodes for DBLP), number of classified graphs, number of features of the node to be classified, number of classes to be predicted, number of continuous features, and number of discrete features.
@@
DBLP 1 26,128 4,057 - 334 4 50 4,565
IMDB 1 11,616 4,278 - 3,066 3 6,132 3,066
MUTAG 188 17.93 - 188 - 2 - -
BA-Shapes 1 700 700 - - 4 - -
Tree-Cycle 1 871 871 - - 2 - -
Tree-Grids 1 1,231 1,231 - - 2 - -
BA-3Motif 3,000 21.92 - 3,000 - 3 - -
@@
We run experiments for DiGNNExplainer on the following datasets in Table 5: DBLP (Fu et al. 2020) is a heterogeneous dataset of a citation network with four node types—Author, Paper, Term, and Conference. Author node features are bag-of-words representations of their paper keywords. Paper nodes have bag-of-words features, Term nodes have continuous features, and Conference nodes have no features. IMDB (Fu et al. 2020) is a heterogeneous dataset for movies with three node types—Movie, Director, and Actor. The Movie node features include bag-of-words representations of their plot keywords. Director and Actor nodes have continuous features. MUTAG (Morris et al. 2020) is a molecule graph classification dataset where each graph can be classified into two classes—mutagenic and non-mutagenic, and each node has a single node type, atom. The synthetic BA-Shapes (Ying et al. 2019) dataset is constructed by randomly attaching house motifs to the nodes of a Barabási–Albert (BA) base graph. All nodes are of a single type. Tree-Cycle (Ying et al. 2019) and Tree-Grids (Ying et al. 2019) are constructed by randomly attaching cycle motifs and n-by-n grid motifs respectively, to the nodes of a balanced binary tree base graph. BA-3Motif (Wang et al. 2021) is a motif graph classification dataset where a house/cycle/grid motif is attached randomly to the nodes of Barabási–Albert(BA) base graphs. The graphs can be classified into any of three types, as per the type of motif attached.
Appendix D Explainers
We briefly describe the implementation of our XGNN and GNNInterpreter baselines for node classification.
XGNN.
In the original implementation of XGNN (Yuan et al. 2020) for MUTAG dataset, explanation graph generation starts by initializing a graph with a single node. Further nodes and edges are added to this graph step-by-step based on the reward used in reinforcement learning. The reward at a given step, is positive if the generated graph at that step is predicted correctly by the trained GCN, otherwise the reward is negative. We extend XGNN in a similar way, starting graph generation with a single-node graph. For node classification, (1) we modify the code to make predictions for node labels and (2) we reduce the threshold value used in the original reward computation in order to obtain connected explanation graphs for our datasets.
GNNInterpreter.
In the original implementation of GNNInterpreter for MUTAG dataset, graph pooling is performed to obtain a graph embedding for each graph. Then a mean embedding is created for graphs of each class, which is used to sample graphs of each class. We extend GNNInterpreter for node classification datasets. (1) We create average embedding of each node class using node pooling. Thus, we can similarly sample explanation graphs of each node class. (2) We fine-tune the hyperparameters for our datasets for each node class to obtain connected explanation graphs.
Appendix E Community Detection [^5] https://python-louvain.readthedocs.io/en/latest/api.html
We use communities or motifs to evaluate if the explanation graphs achieve ground-truth faithfulness. Our communities are similar to ground truth motifs (Ying et al. 2019). Existing techniques (Subramonian 2021; Yu and Gao 2022) use a blackbox model for motif detection. Our community detection technique identifies the building blocks (see Table 8) of a real-world graph dataset by obtaining the largest substructures or communities  (Blondel et al. 2008) in the input graphs sampled from our datasets. This is done similarly in GraphXAI (Agarwal et al. 2023), which uses substructure matching to detect ground-truth in real-world dataset, MUTAG. We generate motifs for synthetic datasets using ShapeGGen  (Agarwal et al. 2023). An explanation achieves ground-truth faithfulness if it contains the community. If the building blocks of the explanation graphs that maximize predictions also achieve a high ground-truth faithfulness score i.e., are same as the building blocks of the real-world data, then the prediction using the trained GNN model represents the real-world behavior (Yuan et al. 2020). In this case, the GNN predictions can be trusted, and our explanation graphs are good explanations.
We can obtain motifs for each class for the real node classification datasets DBLP and IMDB. In DBLP, most of the real graphs (104 out of 120) sampled using ForestFireSampler have all Author nodes of the same class. For IMDB, only some graphs have Movie nodes of the same class (40 out of 143). If a sampled graph has Author nodes of different classes, i.e., classes 1 and 3, we do not use it for motif detection, because the obtained motifs might be common across classes. In this way, we detect motifs that help to distinguish different classes, such that for a DBLP motif of class 0, all author nodes in the real graph belong to class 0. For the datasets BA-Shapes, Tree-Cycle and Tree-Grids, we need to use common motifs for all classes, as we classify the node labels (and not the node classes as in Author) and all sampled graphs have nodes of all labels. For example, in BA-Shapes, a house motif has all four node labels. Table 6 shows motif statistics for common motifs.
Table 6: Motif Statistics: total number of motifs, number of motifs that appear in the best explanation graph of every class, in at least $\frac{1}{2}$ of the classes, and in none of the classes.
Dataset No. of motifs No. of motifs No. of motifs No. of motifs
(every class) ($\frac{1}{2}$ of the classes) (none of the classes)
DBLP 30 13 21 4
IMDB 30 26 28 2
BA-Shapes 30 14 26 4
Tree-Cycle 30 15 28 2
Tree-Grids 30 28 30 0
Appendix F Validity Analysis

### v2

@@ 4.5. Experiment on Feature Selection
In Table 2, we select varying numbers of features (2, 4, 8, 50 for DBLP and 2, 5, 10, 3066 for IMDB) using Variance Threshold. As an alternative, Table 11 in the Appendix reports results using the top $k$ most frequent features as a selection method. Irrespective of the feature selection method, we obtain good explanation graphs, for example, with 50 features using DiTabDDPM on DBLP.
4.6. Ablation Study
We conduct an ablation study to investigate the effect of using multiple DiGress models for different graph sizes versus a fixed-size graph model. We report predictive faithfulness and ground-truth faithfulness scores for three settings: (1) a single diffusion model trained on a fixed graph size (12 for DBLP and 7 for IMDB; Table 4, left), (2) a single diffusion model trained on multiple graph sizes (10–15; Table 4, right), and (3) our main setting, which trains a separate DiGress model for each graph size (10–15; Table 2). Overall, we observe that training one model per graph size yields the best performance. In particular, for DBLP, training a single model on multiple sizes results in a lower ground-truth faithfulness of 0.730, compared to 0.882 in the main experiments.
Table 4. Comparison of predictive faithfulness (PF) and ground-truth faithfulness (GF) for ablation study using a fixed graph size versus using multiple graph sizes by training a single Diffusion model.
Approach Fixed Multiple
PF $\uparrow$ GF $\uparrow$ PF $\uparrow$ GF $\uparrow$
DBLP
DiTabDDPM (50 feat.) 0.999 $\pm$ 0.0 0.717 $\pm$ 0.040 0.999 $\pm$ 0.0 0.730 $\pm$ 0.036
DiTabDDPM (8 discr. feat.) 0.872 $\pm$ 0.102 0.787 $\pm$ 0.037 0.971 $\pm$ 0.030 0.777 $\pm$ 0.043
DiTabDDPM (4 discr. feat.) 0.770 $\pm$ 0.141 0.792 $\pm$ 0.043 0.770 $\pm$ 0.110 0.829 $\pm$ 0.071
DiTabDDPM (2 discr. feat.) 0.925 $\pm$ 0.073 0.735 $\pm$ 0.066 0.938 $\pm$ 0.071 0.812 $\pm$ 0.057
IMDB
DiTabDDPM (3066 feat., all) 0.976 $\pm$ 0.058 0.860 $\pm$ 0.051 0.983 $\pm$ 0.021 0.853 $\pm$ 0.026
DiTabDDPM (10 discr. feat.) 0.967 $\pm$ 0.056 0.863 $\pm$ 0.069 0.972 $\pm$ 0.050 0.853 $\pm$ 0.030
DiTabDDPM (5 discr. feat.) 0.951 $\pm$ 0.087 0.886 $\pm$ 0.022 0.956 $\pm$ 0.065 0.883 $\pm$ 0.026
DiTabDDPM (2 discr. feat.) 0.902 $\pm$ 0.099 0.836 $\pm$ 0.037 0.934 $\pm$ 0.100 0.876 $\pm$ 0.033
4.7. Runtime Analysis
@@
We compare our runtime against the baselines XGNN, GNNInterpreter, and D4Explainer (Table 5). Unlike graph generation techniques that primarily assess scalability with respect to increasing graph size, we evaluate scalability with respect to the size of the original dataset. All measurements report the total time needed for all graph sizes. Compared to D4Explainer, DiGNNExplainer is computationally faster on the large datasets DBLP and IMDB. The runtime of DiGNNExplainer remains approximately constant for both small (Tree-Cycle, 871 nodes) and large (DBLP, 26128 nodes) datasets since our approach samples only small explanation graphs.
Moreover, we break down the runtime of our approach by dataset and by component (Table 6) showing that the dominant cost of DiGNNExplainer is graph generation training. For node classification datasets, we additionally run the Forest Fire Sampler (FFS) to extract training subgraphs for each explanation graph size (e.g., 10–15, i.e., six sizes). For each size, we sample $k=200$ subgraphs, resulting in 1,200 sampled graphs in total. We empirically observe that the runtime grows from the small synthetic datasets to IMDB and DBLP, reflecting higher sampling overhead on larger underlying graphs. Next, we train DiGress on these sampled graphs. The Graph Generation Training (GGT) and Graph Generation After Training (GGAT) steps scale linearly with the number of sampled and generated graphs, respectively, as well as with the number of diffusion steps. Additionally, we train a model for node feature generation, where the runtimes for Node Feature Generation Training/After Training (NFGT/NFGAT) are only incurred for datasets with generated node attributes (here: DBLP and IMDB) and depend on the feature dimensionality and model hyperparameters. The Post-Processing (PP) step checks whether the generated graph is consistent with the metagraph. Although graph matching can be computationally hard in the worst case, this step is fast in practice for our small explanation graphs, as confirmed by our empirical runtimes. Finally, the total runtime for Graph Selection (GS), including the required GNN predictions scales linearly with the number of graphs and with the dimensionality of the node features.
@@
Table 5. Runtime (in secs.) comparisons for all graph sizes with baseline approaches.
@@
XGNN 1102 363 27 351 119 155 79
GNNInterpreter 79 102 5 143 73 73 13
D4Explainer 39175 19406 658 5703 659 822 801
DiGNNExplainer 5875 4702 6910 4637 5192 5501 1992
Table 6. Breakdown of runtime (in secs.) for each component (Forest Fire Sampler (FFS), Graph Generation Training (GGT), Graph Generation After Training (GGAT), Node Feature Generation Training (NFGT), Node Feature Generation After Training (NFGAT), Postprocessing ($P$), Graph selection (GS)) of DiGNNExplainer for all datasets.
FFS GGT GGAT NFGT NFGAT $P$ GS
DBLP 7.92 5297.4 255.5 252.10 41.01 20.48 0.72
IMDB 5.08 4054.8 192.3 108.73 36.46 304.57 0.61
MUTAG 6699.4 210.2 0.83
BA-Shapes 3.80 4380.5 252.6 0.69
Tree-Cycle 3.65 4926.1 253.8 0.91
Tree-Grids 3.75 5238.4 258.5 0.79
BA-3Motif 1935.6 49.8 1.49
4.8. Case Study
In this section, we present case study examples of explanation graphs with node features. For each class, we visualize the explanation graph that maximizes the GNN’s predicted probability for that class and corresponding feature distributions. Concretely, we show the feature vector of the author node with the highest probability and the continuous distribution of all term nodes.
Figures 4, 4, and 4 show the explanation graphs and node features for the DiTabDDPM (50 discr. feat.) approach for DBLP. The size of the explanation graphs is between 12 and 15. All author classes are predicted with a probability of 1.0. The ground-truth faithfulness obtained for author classes is 0.8275. Figure 4 shows the distribution of the bag-of-words features (0.0/1.0) of the author node predicted with maximum probability for each class. The term nodes have continuous features (Figure 4) with a mean between 0.04 and 0.11 and a standard deviation between 0.60 and 1.22. The explanations for all approaches and all datasets can be found in our evaluation results. [^5] https://github.com/ds-jrg/DiGNNExplainer/tree/main/evaluation
Nodes 12
Prob. 1.0
(a) class 0
Nodes 15
Prob. 1.0
(b) class 1
Nodes 14
Prob. 1.0
(c) class 2
Nodes 14
Prob. 1.0
(d) class 3
Figure 2. Explanation graphs for each class of author nodes of the DBLP dataset that maximizes the prediction of the class. Node colors indicate types: paper (blue), author (red), term (dark green), and conference (orange).
(e) class 0
(f) class 1
(g) class 2
(h) class 3
Figure 3. Frequency distribution plots visualizing the synthetic bag-of-words node feature vector of the author node of the explanation graph that is classified with maximum probability for a class.
(a) class 0
(b) class 1
(c) class 2
(d) class 3
Figure 4. Density plots of the continuous term-node feature values in the explanation graphs. For each class, we show the feature-value distribution of the term nodes in the explanation graph that maximizes the predicted probability for that class.
5. Discussion
@@
We discuss the following improvements and extensions to our approach: (1) Our approach separates the process of generating explanation graphs into two distinct steps: graph structure generation and node feature generation. This decoupling allows discrete and continuous features to be learned in their appropriate spaces, which performs better than joint continuous-space modeling. However, it does not capture advanced feature-structure dependence. We explored modeling such dependencies in DiTabDDPM and TabDDPM, but observed reduced feature realism, as quantified by cosine similarity. Therefore, we present our main results without considering advanced feature-structure dependence. In our approach, structural information is retained through node types, and feature-structure variation is captured via separate feature models per node type. (2) We obtain consistent runtime for both small and large datasets, making our approach applicable to graphs of varying sizes. Runtime could be further reduced in future work by using multiple GPUs simultaneously, as in recent DiGress implementations. (3) Training separate models for different graph sizes improves our explanations, but entails a trade-off between faithfulness and computational time. A faster alternative is to train a single diffusion model on multiple graph sizes. However, our ablation study (Table 4) shows that this yields weaker evaluation results and generates a higher number of disconnected graphs. (4) Our approach is easily extendable to real-world heterogeneous directed graphs, since our GNNs (GraphSAGE and HAN) can be trained on directed graphs. Moreover, DiGNNExplainer, which builds upon DiGress, is able to support directed graphs with only minor modifications, i.e., by modifying the way noise is added to the adjacency matrix. (5) We currently obtain one explanation per class, but our implementation also supports multiple explanation graphs per class, i.e., the top $k$ explanations ranked by softmax probabilities.
6. Conclusion
@@
DiGNNExplainer is the first method to explain heterogeneous GNNs on a model-level on real-world datasets using explanation graphs with actual node features. Our experiments show that incorporating actual node features significantly improves explanation faithfulness: DiGNNExplainer achieves higher predictive and ground-truth faithfulness than state-of-the-art baselines, while producing explanation graphs whose statistical properties closely match those of real-world graphs. This is enabled by (1) node feature generation via our novel discrete diffusion model, DiTabDDPM, which enables the synthesis of realistic discrete features; and (2) our discrete diffusion-based approach for modeling heterogeneous graphs, supporting the generation of structurally realistic explanation graphs. Beyond the citation and movie datasets used in our experiments, DiGNNExplainer is broadly applicable to any domain with heterogeneous graph data and structured node features, such as healthcare, finance, or chemistry. Further runtime optimization is left for future work.
Acknowledgements.
This work has been supported by the European Union and the Ministry of Economic Affairs, Industry, Climate Action and Energy of the State of North Rhine-Westphalia (MWIKE NRW) under the grant no EFRE-20801043.
@@ Appendix A Diffusion Models
Figure 5. Diffusion steps for generation of synthetic graph and discrete node features.
A.1. DiGress
DiGress (Vignac et al., 2023) is a discrete diffusion approach for graph generation, which adds noise through discrete time steps based on a Markov process. The noise model is a Markov process over a discrete state-space. In a discrete state-space, each node and edge transitions to a different state with the addition of noise at every time step. A state refers to a node type or an edge type. Following each transition, the node or edge changes to another type. The probabilities of transition from one state to another are represented as transition matrices for nodes and for edges. In DiGress, the next step $G_{t}$ (Figure 5) is obtained by the product of the transition matrix, $Q_{t}$ at step $t$ and the graph at the previous step, $G_{t-1}$. The transition probability, $q$ at $G_{t}$ is given as:
(1) $q(G_{t}|G_{t-1})=G_{t-1}Q_{t}$
@@ A.1. DiGress
where node loss, $L_{Nodes}$ and edge loss, $L_{Edges}$ represent the cross-entropy loss between the true node and edge, and predicted probabilities of each node and edge, and $\lambda\in\mathbb{R}^{+}$ assigns relative importance to nodes and edges. The graph loss is computed as a combined loss of individual node loss and edge loss. The loss of each node and edge is measured using cross-entropy. DiGress models node features along with the graph structure. The node features are one-hot encoding of node types. Extra features are added to nodes, including structural information that encodes the presence of cycles, and spectral features for positional encodings needed for a transformer model. These extra features are added at each diffusion step, as the message passing networks are unable to capture the presence of cycles. DiGress performs well at generating small heterogeneous graphs with the same node size as the input graph. Discrete node types and edge types are sampled from the complete noise distributions. The samples are input to a denoising network, which uses self-attention to predict the noise and outputs the distributions for node types and edge types for each previous step.
A.2. TabDDPM
TabDDPM (Kotelnikov et al., 2023) extends the denoising diffusion probabilistic (DDPM) approach used for images to synthesize tabular data. In TabDDPM, an input feature vector, $x=[x_{num},x_{cat_{1}}\dots x_{cat_{C}}]$ consists of both numeric features, $x_{num}$ and categorical features, $x_{cat_{i}}$ with $i\in C$ categories. The categorical features are represented as one-hot encoding of categories. TabDDPM models categorical and binary features using multinomial diffusion and numerical features using continuous noise. The forward process adds Gaussian noise using a Markov process, while the reverse process uses a denoising neural network consisting of MLP layers with SiLU non-linear activation. The denoising network is an MLP architecture of the form (Gorishniy et al., 2021):
(2) $\mathit{MLP}(x)=Linear(\mathit{MLPBlock}(\dots(\mathit{MLPBlock}(x))))$
@@ Appendix B Additional Details for DiGNNExplainer
B.1. DiTabDDPM - Denoising network
@@
B.2. Node Feature Reconstruction Loss
@@
$L_{recon}=\mathbb{E}\left[\log{p_{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{t})\right]$
B.3. Postprocessing
Two nodes $\tau_{1},\tau_{2}$ in the metagraph are connected if and only if there exist two nodes $v_{1}$ and $v_{2}$ in the graph such that $v_{1}$ has type $\tau_{1}$ and $v_{2}$ has type $\tau_{2}$. The generated graph is consistent if and only if the metagraph of the generated graph is a subgraph of the metagraph of the dataset.
B.4. Explanation Graph Selection
@@ Appendix C Datasets
Table 7. Dataset statistics: number of graphs, average number of nodes per graph, number of classified nodes (e.g., author nodes for DBLP), number of classified graphs, number of features of the node to be classified, number of classes to be predicted, number of continuous features, and number of discrete features.
@@
DBLP 1 26,128 4,057 334 4 50 4,565
IMDB 1 11,616 4,278 3,066 3 6,132 3,066
MUTAG 188 17 188 2
BA-Shapes 1 700 700 4
Tree-Cycle 1 871 871 2
Tree-Grids 1 1,231 1,231 2
BA-3Motif 3,000 21 3,000 3
@@
We run experiments for DiGNNExplainer on the following datasets in Table 7: DBLP (Fu et al., 2020) is a heterogeneous dataset of a citation network with four node types—author, paper, term, and conference. Author node features are bag-of-words representations of their paper keywords. Paper nodes have bag-of-words features, term nodes have continuous features, and conference nodes have no features. IMDB (Fu et al., 2020) is a heterogeneous dataset for movies with three node types—movie, director, and actor. The movie node features include bag-of-words representations of their plot keywords. Director and actor nodes have continuous features. MUTAG (Morris et al., 2020) is a molecule graph classification dataset where each graph can be classified into two classes—mutagenic and non-mutagenic, and each node has a single node type, atom. The synthetic BA-Shapes (Ying et al., 2019) dataset is constructed by randomly attaching house motifs to the nodes of a Barabási–Albert (BA) base graph. All nodes are of a single type. Tree-Cycle (Ying et al., 2019) and Tree-Grids (Ying et al., 2019) are constructed by randomly attaching cycle motifs and n-by-n grid motifs, respectively, to the nodes of a balanced binary tree base graph. BA-3Motif (Wang et al., 2021) is a motif graph classification dataset where a house/cycle/grid motif is attached randomly to the nodes of Barabási–Albert (BA) base graphs. The graphs can be classified into any of three types, as per the type of motif attached.
Appendix D Validity Analysis